OpenStudy (anonymous):

without using a calculator find the exact value of cos^-1 (cos 17pi/5)

OpenStudy (jdoe0001):

hmmm, what does \(\bf cos^{-1}(whatever)\) means?

OpenStudy (anonymous):

the inverse

OpenStudy (debbieg):

Yes, now think about that. The inverse cosine gives the ANGLE that has, as its cosine value, the INPUT of the inverse function, right?

OpenStudy (jdoe0001):

the inverse? well, ahemm... \(\bf cos^{-1}(whatever) \) means, what's the angle WHOSE COSINE is "whatever"

OpenStudy (jdoe0001):

so, what would be \(\bf cos(whoever) \)?

OpenStudy (jdoe0001):

or I should say what would be \(\bf cos(someangle)\)

OpenStudy (jdoe0001):

the cosine function is just a ratio, a number, a value you give it an angle, it gives you the ratio value, some number \(cos^{-1}\) takes THAT NUMBER and spits out the angle it came from

OpenStudy (anonymous):

17pi/5, but aren't there restrictions on the inverse of cosine like it only is in the first two quadrants from 0 to pi? so then wouldn't 17pi/5 not work?

OpenStudy (jdoe0001):

yes, when you have the actual value, yes the range restriction will show up when you know the source, then you don't have to do anything with the restrictions

OpenStudy (debbieg):

Yes, now you have the find the function that is IN the output of the inverse cosine function, that wold have the SAME cosine as 17pi/5

OpenStudy (jdoe0001):

\(\bf cos(2\pi)=1, cos(50\pi)=1\\ cos^{-1}(1) = 0\\ \textit{but 0 is a coterminal with }50\pi\\ cos^{-1}(cos(50\pi)) = 50\pi\)

OpenStudy (debbieg):

17pi/5 - 2pi=7pi/5 so is in the 3rd quadrant, so not in the range of inverse cosine (which is limited to 1st and 2nd quadrant). But we know that it has reference angle of 7pi/5-pi=7pi/5-5pi/5=2pi/5. Since in Q3 it has negative cosine. So the angle between 0 and pi that has the same cosine is.....?

OpenStudy (debbieg):

Sorry but I will disagree with @jdoe0001 on this point. The output range for the function of 0 to 2pi, 50pi is not in that range, so can't be the value of the inverse cosine. I have given my students problems designed exactly like this to test their understanding of this. cos(50pi) is a NUMBER that is in the domain of inverse cosine, but it gives the result =0, not =50pi, because 50pi is not in the range of inverse cosine.

OpenStudy (anonymous):

@DebbieG ok i understand how to get the 7pi/5, but then how do you know to subtract pi?

OpenStudy (debbieg):

The angle 7pi/5 is in quadrant 3, right? Do you know what I mean by "reference angle"? (Different texts may have different terminology.) That is the acute, positive angle between the terminal side of any (non-quadrantal) angle and the x-axis. So, there is one angle in each quadrant that has a given (acute) reference angle, and each of those 4 angles has the same value for cosine, EXCEPT POSSIBLY for the sign of it. E.g., COS(pi/3)=1/2, COS(2pi/3)=-1/2, COS(4pi/3)=-1/2, COS(5pi/3)=1/2. All have the same reference angle. So now think about 7pi/5. where is it? it is 2pi/5 "past" pi (7pi/5-pi=2pi/5). So that's the reference angle. So COS(7pi/5)=-COS(2pi/5), and COS(7pi/5)=COS(3pi/5).... do you see why? (Think about which quadrant each is in, and what the reference angle is.)

OpenStudy (anonymous):

@DebbieG oh ok i was confused i thought the answer you were saying was 2pi/5, but that's the reference angle and the actual answer is 3pi/5 thank you so much for your help!!!

OpenStudy (debbieg):

yes indeed, sorry for the confusion.... :)